OpenStudy (anonymous):

Factor the following polynomial: x2 – 3x – 40

OpenStudy (anonymous):

8 and 5.

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

(x – 8)(x + 5)

OpenStudy (stamp):

\[x^2-3x-40\] c = 8 * 5, 10 * 4, 20 * 2, 40 * 1 What +/- combination obtains b = - 3? +/- 8 +/5 works. -8 +5 = -3 (x - 8) (x + 5)

OpenStudy (anonymous):

thank you!

OpenStudy (stamp):

Check (x-8)(x+5) x(x+5)-8(x+5) xx + 5x -8x- 40 xx -3x -40 x^2 - 3x - 40